(2) Let r1 1 and -(-) 1 (n+1)2 = I+ur (a) Show that lim,, T, exists. (b) Prove that z, #t1 by induction and find the limit.
Answer #1

りとり nin 6 mal, and mnt =(1- M712) n So, mut = (1+ moti) (1-he) in = n2 x more nn = Menn New by induction. We prove that N Abn Now, by induction we have to prove that ng = htt ; tips for n=s, the statement is true. let us assume that the statement i